Hale Corporation is comparing two different capital structures, an all-equity plan (Plan I) and a levered plan (Plan II). UndHale Corporation is comparing two different capital structures, an all-equity plan (Plan I) and a levered plan (Plan II). Under Plan I, the company would have 200,000 shares of stock outstanding. Under Plan II, there would be 150,000 shares of stock outstanding and $2.2 million in debt outstanding. The interest rate on the debt is 5 percent and there are no taxes.
Answer #1

a.

Plan I Plan II
EBIT 350,000 350,000
less:interest on debt (2.2 m *5%) nil (110,000)
EBT 350,000 240,000
EPS (EBT / number of shares) $1.75 $1.60
EPS
Plan I 1.75
Plan II 1.60

b.

Plan I Plan II
EBIT 600,000 600,000
less:interest nil (110,000)
EBT 600,000 490,000
EPS (EBT / number of shares) $3 $3.27
EPS
Plan I 3
plan II 3.27

c.

working:

at break even EBIT = EPS is same for both the plans.

let x be the EBIT.

EPS of plan I = (x - interest) / number of shares

=> (x-0)/ 200,000

EPS of plan II = (x -110,000) / 150,000

so at break even point both EPS are equal

=> x/200,000 = (x-110,000) /150,000

=>0.75x = x-110,000

=>x=110,000/0.25

=>440,000.